OpenStudy (anonymous):

in the xy-coordinate system,if (-3,B) and (A-3,B-3)are two points on the line defined by x+3y=12 the A=?

OpenStudy (anonymous):

No can do Ma'am, I am tired and going to sleep.

OpenStudy (anonymous):

@lgbasallote

OpenStudy (anonymous):

@.Sam. can u solve it?

OpenStudy (anonymous):

@mikala1

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

@calculusfunctions

OpenStudy (anonymous):

CAN YOU HELP HER

OpenStudy (anonymous):

plz help

OpenStudy (raden):

first, because (-3,B) on the line of x+3y=12, so just plug x=-3 and y=B to this line

OpenStudy (raden):

it means that -3+3(B) = 12, solve for B

OpenStudy (calculusfunctions):

Alright so we have (-3, B) and (A - 3, B - 3) are on the line x + 3y = 12. Correct?

OpenStudy (anonymous):

after that?

OpenStudy (calculusfunctions):

@RadEn I thought you knew better than to give out the answers?

OpenStudy (anonymous):

i got it the answer is 9

OpenStudy (raden):

sure, i want @girl to solve it

OpenStudy (anonymous):

what is the product of the distinct prime factors of 150?

OpenStudy (raden):

hmmmm... is it the other question ?

OpenStudy (calculusfunctions):

@girlish do you want to learn or do you just want answers, because if it's just answers you're looking for then please tell me so that I don't waste my time.

OpenStudy (anonymous):

yes @RadEn i want to learn as in solution and answer both @calculusfunctions

OpenStudy (calculusfunctions):

NO!!! I do not give out solutions or answers! I will teach you to arrive at the correct solution. Please tell me quickly because I don't have a lot of time because I have to go. I came to help you because it was a request but I don't have time to play games.

OpenStudy (anonymous):

i want to learn plz help

OpenStudy (anonymous):

m not playing games here

OpenStudy (raden):

well, @calculusfunctions was off line. can u solve -3+3B = 12 ?

OpenStudy (anonymous):

he left because u took to long sorry

OpenStudy (anonymous):

yes ive solved

OpenStudy (anonymous):

what is the product of the distinct prime factors of 150?

OpenStudy (raden):

what;s value of B do u get ? the 2nd is eassy but i want u finish number one, first

OpenStudy (anonymous):

i got 5

OpenStudy (raden):

ok, good.... next we are going to the 2nd step. line x+3y=12 passes point (A-3,B-3), it means that : A-3+3(B-3) = 12, plug B=5 to this equation and solve for A

OpenStudy (anonymous):

yes i know, A wud be 9.Right?

OpenStudy (raden):

ok, the first was solved :) u are right

OpenStudy (anonymous):

now come to the second one please :)

OpenStudy (raden):

where is it ?

OpenStudy (anonymous):

what is the product of the distinct prime factors of 150?

OpenStudy (anonymous):

this is the second one

OpenStudy (raden):

okeeeey :) look this ilustration